OpenStudy (anonymous):

A company's monthly profit, P, from a product is given by P = −x2 + 105x − 1050, where x is the price of the product in dollars. What is the lowest price of the product, in dollars, that gives a monthly profit of $1,550? (Do not put a $ in your answer.)

Parth (parthkohli):

Replace \(P = 1,550\). \(1550 = -x^2 + 105x - 1050\) A start is subtracting 1550 from both sides.

OpenStudy (anonymous):

-500

OpenStudy (hba):

1550 = -x^2 + 105x - 1050 -x^2 + 105x - 2600 =0 x= 40 or x=65 lowest price =40

OpenStudy (hba):

bbtw -1050-1550=-2600 not -500

Parth (parthkohli):

@hba Please do not answer questions directly without letting the asker respond. http://openstudy.com/code-of-conduct

OpenStudy (hba):

@ParthKohli Rather You should ask @MeganLewis To respond .

Parth (parthkohli):

Well, you are not allowed to give answers.

OpenStudy (hba):

You are wrong

OpenStudy (hba):

you are not allowed to give answers directly @ParthKohli

Parth (parthkohli):

No, you are not allowed to give answers at all. Most of the people are not following that rule.

OpenStudy (hba):

You are wrong

Parth (parthkohli):

You may give hints, but not answers.

OpenStudy (hba):

There is no such rule

Parth (parthkohli):

Actually, there is. @Callisto @asnaseer Can you please explain?

OpenStudy (hba):

@Callisto Yes Please

OpenStudy (anonymous):

the code states "dont post only answers" so its possible to give answer as long as you explain it

OpenStudy (hba):

^^ Right

Parth (parthkohli):

"Guide the askers to a solution".

OpenStudy (hba):

^^ You Are Wrong Parth

Parth (parthkohli):

We must 'guide' the asker. Guiding is different than giving answers.

OpenStudy (hba):

I Gave the Whole Solution So I Cant be Given a warning

Parth (parthkohli):

Let me explain it to you. There are some people known as 'answer fiends' who are just concerned with the answers, NOT the solutions. So even if you give the solution, it does no good as the answer fiends will see only the last step, i.e the answer. The site prevents some legal issues by passing rules such as this one.

OpenStudy (asnaseer):

ParthKohli is right in this case hba. You did all the work up there and left nothing for the asker of this question to do. It is better to guide someone to an answer, letting them take part in the process. That way, the asker learns far more.

OpenStudy (asnaseer):

I am sure your intention was sincere hba - but I hope you understand the difference

OpenStudy (hba):

Ok No Problemo

OpenStudy (asnaseer):

Thanks for understanding hba - much appreciated. :)

OpenStudy (hba):

Your Welcome

OpenStudy (anonymous):

@hba Thanks! I get it now.